OpenStudy (anonymous):

Please help! Will fan and medal! a. How many real solutions does the equation 0.2x^5 -2x^3 +1.8x +k =0 have when k = 0? b. How many real solutions does it have when k=1? c. Is there a value of k for which the equation has just one real solution? d. Is there a value of k for which the equation has no real solution?

OpenStudy (anonymous):

Here's the equation again in case the one I typed out in the question is difficult to read: \[0.2x^5-2x^3+1.8x+k=0\]

OpenStudy (imstuck):

Do you know Descartes' Rule of Signs?

OpenStudy (anonymous):

I do not.

OpenStudy (ikram002p):

0.2x^5 -2x^3 +1.8x +k =0 thus 0.2x^5 -2x^3 +1.8x =0 x(0.2x^4-2x^2+1.8 )=0

OpenStudy (anonymous):

So what do I do with the equation you've given me to answer the questions above?

OpenStudy (anonymous):

@ikram002p